a stone is thrown horizontally from the roof of a building 15 m high with an initial speed of 5 m/s. How long will it take before the ball hits the ground?


1
Expert's answer
2022-06-01T12:48:54-0400

The time of motion is equal to the time of free falling:

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2*15}{9.8}}=1.75\:\rm s"